OpenStudy (anonymous):

I need someone to check my work, please: (a2/3)(a5/4) (the fractions are exponents) I came up with: a5/6 (the fraction is an exponent

OpenStudy (precal):

bases are the same add the exponents (2/3) + (5/4) (8/12)+(15/12) 23/12 is the power

OpenStudy (anonymous):

2/3 + 5/4 - what precal said

OpenStudy (precal):

recall x(x) is x^2 because 1+1 = 2 x(x^2)=x^3 becasue 1+2=3

OpenStudy (anonymous):

okay so a7/7, but do I take it further?

OpenStudy (anonymous):

no the answer is a^(23/12) (its best to use ^ to indicate and exponent)

OpenStudy (anonymous):

Got you! Thanks so much for the help!

OpenStudy (anonymous):

* indicate an exponent

OpenStudy (anonymous):

yw